The difference of two numbers is 36. The larger is 6 less than 4 times the smaller. What are the two numbers?
Answer by algebrahouse.com(1079)

"The difference of two numbers is 36. The larger is 6 less than 4 times the smaller. What are the two numbers?"

x = smaller
4x - 6 = larger {larger is six less than four times smaller}

4x - 6 - x = 36 {difference of two numbers is 36
3x - 6 = 36 {combined like terms}
3x = 42 {added 6 to both sides}
x = 14 {divided both sides by 3}
4x - 6 = 50 {substituted 14, in for x, into 4x - 6}

smaller number is 14
larger number is 50
www.algebrahouse.com